# Constraints on Null Energy Condition Violation from Advanced LIGO and Advanced Virgo’s First Three Observing Runs

### v1

@@ 4 Results and discussions
To assess the presence of the SGWB signal originating from the NEC-violating phase, we calculate the Bayes factor $\mathcal{B}_{\text{NOISE}}^{\text{GW}}=0.53$. This value suggests that there is no significant evidence of such a signal in the current data. Consequently, we impose an upper limit on the power spectrum amplitude during the second inflationary stage, $P_{T,2}$. At a $95\%$ confidence level, we constrain $P_{T,2}\lesssim 0.15$. This bound may be applied to specific inflationary models that incorporate NEC violation, such as those featuring non-canonical kinetic terms Creminelli:2006xe , thereby restricting their viable parameter space.
Notably, our derived upper limit on $P_{T,2}$ is consistent with the constraints obtained from PTA observations Ye:2023tpz . This consistency highlights the complementary nature of PTA and Advanced LIGO-Virgo observations in probing the physics of the early Universe. PTAs are sensitive to GWs in the nanohertz frequency range, while Advanced LIGO-Virgo explores higher frequencies between 10 and 200 Hz. The agreement between the constraints obtained from these different frequency regimes strengthens the case for the NEC-violating inflationary scenario as a plausible explanation for the PTA signal. As future GW experiments with improved sensitivities come online, we anticipate enhanced capabilities in testing inflationary models and further constraining their parameters.

### v2

@@ 4 Results and discussions
To assess the presence of the SGWB signal originating from the NEC-violating phase, we calculate the Bayes factor $\mathcal{B}_{\text{NOISE}}^{\text{GW}}=0.53$. This value suggests that there is no significant evidence of such a signal in the current data. Consequently, we impose an upper limit on the power spectrum amplitude during the second inflationary stage, $P_{T,2}$. At a $95\%$ confidence level, we constrain $P_{T,2}\lesssim 0.15$. This bound may be applied to specific inflationary models that incorporate NEC violation, such as those featuring non-canonical kinetic terms [51], thereby restricting their viable parameter space.
Notably, our derived upper limit on $P_{T,2}$ is consistent with the constraints obtained from PTA observations [42]. This consistency highlights the complementary nature of PTA and Advanced LIGO-Virgo observations in probing the physics of the early Universe. PTAs are sensitive to GWs in the nanohertz frequency range, while Advanced LIGO-Virgo explores higher frequencies between 10 and 200 Hz. The agreement between the constraints obtained from these different frequency regimes strengthens the case for the NEC-violating inflationary scenario as a plausible explanation for the PTA signal. As future GW experiments with improved sensitivities come online, we anticipate enhanced capabilities in testing inflationary models and further constraining their parameters.